# A Quantum of Learning: Using Quaternion Algebra to Model Learning on Quantum Devices

### v1

@@ II Quaternion Algebra
Most notably, it follows from (1) that multiplication is not an commutative affair in the quaternion domain. [^1] For example, form (1) note that $\jmath=\kappa\imath$, and thus, $\jmath\imath=\kappa\imath\imath=-\kappa\neq\imath\jmath$. The conjugate and norm of $q\in\mathbb{H}$ are respectively given by
@@ III Formulating Quantum Computation with Quaternion Algebra
Although the two complex-valued variables $\alpha$ and $\beta$ might suggest four degrees of freedom for $|\psi\rangle$, the unit sum condition in (5) limits the degrees of freedom to two. Thus, the state of $|\psi\rangle$ can be described via a pure imaginary quaternion as demonstrated in Fig. 2 where
@@
Proof of Proposition 1: Consider the qubit $|\psi_{\text{out}}\rangle$ and $|\psi_{\text{in}}\rangle$ related via quantum gate $\mathbf{U}$, so that $|\psi_{\text{out}}\rangle=\mathbf{U}|\psi_{\text{out}}\rangle$, where $\mathbf{U}\in\mathbb{C}^{2\times 2}$ is a unitary [^2] A unitary transform preserves norm of the input state, ensuring that amplitudes of the input and output states remain one. transform, $\langle\psi_{\text{out}}|\psi_{\text{out}}\rangle=1$, and $\langle\psi_{\text{in}}|\psi_{\text{in}}\rangle=1$ to satisfy the unit sum condition in (5). As a result, quaternions $q_{\text{out}}$ and $q_{\text{in}}$ modelling $|\psi_{\text{out}}\rangle$ and $|\psi_{\text{in}}\rangle$, will be pure imaginary quaternion-valued variables with unit norm. It follows from geometric algebra that $q_{\text{in}}$ and $q_{\text{out}}$ denote a plane in the quaternion imaginary sub-space [28, 29]. In this setting, $\eta=\Im\{q_{\text{in}}q_{\text{out}}\}$ is normal to the plane that contains $q_{\text{out}}$ and $q_{\text{in}}$ [29]. Therefore, if $\theta$ denotes the angle between $q_{\text{out}}$ and $q_{\text{in}}$ in the plane that contains them, the quaternion involution in (2) formulates the relation between $q_{\text{in}}$ and $q_{\text{out}}$ modelling the operation of quantum gate $\mathbf{U}$ in the quaternion domain. This operation is linear in the augmented quaternion space [^3] Also referred to as widely-linear in the literature, see [21, 26]. as expressed in (4). Since in (4) the input and output have the same amplitude, the transform is unitary. Now, consider the $m$-qubits system, the state of which in representable via $\mathbf{q}=\frac{1}{\sqrt{m}}\begin{bmatrix}q_{1},\ldots,q_{m}\end{bmatrix}^{\mathsf{T}}$, where $q_{i}$ represents state of the $i^{\text{th}}$ qubit. Due to the nature of quantum particles, two class of operations are possible; i) a rotation of each qubit, ii) summation and re-normalisation of subset of qubits or their rotated versions [11, 17, 16], both of which are linear transformations in the quaternion augmented space [21]. Therefore, all possible computations on $m$-qubit systems are linear unitary transforms in the augmented quaternion space.
@@ IV Quantum Learning
Let us consider the data set $\mathcal{D}=\{\mathcal{D}_{n}:n\in\mathbb{N}\}$ with sub-set $\mathcal{D}_{n}\sim\mathsf{P}\left(\mathcal{D}_{n}|\zeta\right)$ where $\zeta$ denotes a hypothesis variable. The aim is to find the mapping $\mathsf{f}:\mathsf{P}\left(\mathcal{D}_{n}|\zeta\right)\rightarrow\mathsf{P}\left(\zeta|\mathcal{D}_{n}\right)$, that is, a mapping between probability distribution of a data set and that of the possible hypothesis generating that data set. This is a class of problem with application in Bayesian learning [30], engineering [31], and sequential filtering [7, 8]. In order to formulate the quantum learning problem, the probability distribution $\mathsf{P}\left(\mathcal{D}_{n}|\zeta\right)$ is considered to be encoded as an multi-qubit system denoted with $\mathbf{x}_{n}$, whereas probability distributions $\mathsf{P}\left(\zeta|\mathcal{D}_{n}\right)$ is encoded and denoted with $\mathbf{y}_{n}$.
@@ V Numerical Example
Figure 3: The cost, $10\log\left(\frac{1}{8}\sum^{8}_{s=1}\mathsf{d}\left(\hat{d}_{s,n},d_{s,n}\right)\right)$, is shown for $10^{3}$ independent trials of the derived algorithm. The cost across all trial lie within the light red region, while the solid red line showing the average over all trials.

### v2

@@ II Quaternion Algebra
Most notably, it follows from (1) that multiplication is not a commutative affair in the quaternion domain. [^1] For example, form (1) note that $\jmath=\kappa\imath$, and thus, $\jmath\imath=\kappa\imath\imath=-\kappa\neq\imath\jmath$. The conjugate and norm of $q\in\mathbb{H}$ are respectively given by
@@ III Formulating Quantum Computation with Quaternion Algebra
Although the two complex-valued variables $\alpha$ and $\beta$ might suggest four degrees of freedom for $|\psi\rangle$, the unit sum condition in (5) and Remark 2, limit the degrees of freedom to two. Thus, the state of $|\psi\rangle$ can be described via a pure unit imaginary quaternion as demonstrated in Fig. 2 where
@@
Proof of Proposition 1: Consider the qubit $|\psi_{\text{out}}\rangle$ and $|\psi_{\text{in}}\rangle$ related via quantum gate $\mathbf{U}$, so that $|\psi_{\text{out}}\rangle=\mathbf{U}|\psi_{\text{in}}\rangle$, where $\mathbf{U}\in\mathbb{C}^{2\times 2}$ is a unitary [^2] A unitary transform preserves norm of the input state, ensuring that amplitudes of the input and output states remain one. transform, $\langle\psi_{\text{out}}|\psi_{\text{out}}\rangle=1$, and $\langle\psi_{\text{in}}|\psi_{\text{in}}\rangle=1$ to satisfy the unit sum condition in (5). As a result, quaternions $q_{\text{out}}$ and $q_{\text{in}}$ modelling $|\psi_{\text{out}}\rangle$ and $|\psi_{\text{in}}\rangle$, will be pure imaginary quaternion-valued variables with unit norm. It follows from geometric algebra that $q_{\text{in}}$ and $q_{\text{out}}$ denote a plane in the quaternion imaginary sub-space [28, 29]. In this setting, $\eta=\Im\{q_{\text{in}}q_{\text{out}}\}$ is normal to the plane that contains $q_{\text{out}}$ and $q_{\text{in}}$ [29]. Therefore, if $\theta$ denotes the angle between $q_{\text{out}}$ and $q_{\text{in}}$ in the plane that contains them, the quaternion involution in (2) formulates the relation between $q_{\text{in}}$ and $q_{\text{out}}$ modelling the operation of quantum gate $\mathbf{U}$ in the quaternion domain. This operation is linear in the augmented quaternion space [^3] Also referred to as widely-linear in the literature, see [21, 26]. as expressed in (4). Since in (4) the input and output have the same amplitude, the transform is unitary. Now, consider an $m$-qubits system, the state of which in representable via $\mathbf{q}=\frac{1}{\sqrt{m}}\begin{bmatrix}q_{1},\ldots,q_{m}\end{bmatrix}^{\mathsf{T}}$, where $q_{i}$ represents state of the $i^{\text{th}}$ qubit. Due to the nature of quantum particles, two class of operations are possible; i) a rotation of each qubit, ii) summation and re-normalisation of subset of qubits or their rotated versions [11, 17, 16], both of which are linear transformations in the quaternion augmented space [21]. Therefore, all possible computations on $m$-qubit systems are linear unitary transforms in the augmented quaternion space.
@@ IV Quantum Learning
Let us consider the data set $\mathcal{D}=\{\mathcal{D}_{n}:n\in\mathbb{N}\}$ with sub-set $\mathcal{D}_{n}\sim\mathsf{P}\left(\mathcal{D}_{n}|\zeta\right)$ where $\zeta$ denotes a hypothesis variable. The aim is to find the mapping $\mathsf{f}:\mathsf{P}\left(\mathcal{D}_{n}|\zeta\right)\rightarrow\mathsf{P}\left(\zeta|\mathcal{D}_{n}\right)$, that is, a mapping between probability distribution of a data set and that of the possible hypothesis generating that data set. This is a class of problems with application in Bayesian learning [30], engineering [31], and sequential filtering [7, 8]. In order to formulate the quantum learning problem, the probability distribution $\mathsf{P}\left(\mathcal{D}_{n}|\zeta\right)$ is considered to be encoded as a multi-qubit system denoted with $\mathbf{x}_{n}$, whereas probability distributions $\mathsf{P}\left(\zeta|\mathcal{D}_{n}\right)$ is encoded and denoted with $\mathbf{y}_{n}$.
@@ V Numerical Example
Figure 3: The cost, $10\log\left(\frac{1}{8}\sum^{8}_{s=1}\mathsf{d}\left(\hat{d}_{s,n},d_{s,n}\right)\right)$, is shown for $10^{3}$ independent trials of the derived algorithm. The cost across all trials lie within the light red region, while the solid red line showing the average over all trials.